An investment advisor believes that there is a 28% chance of making money by investing in a specific stock. If the stock makes money, then there is a 43% chance that among those making money, they would also get a dividend. Find the probability that the investor makes money and receive a dividend.

43%
15%
28%
12%
Answer #1

Probability that an investment adviser makes money, P(makes money) = 0.28

Probability that if the stock makes money then they would also get a dividend, P(Get dividend |make money) = 0.43

Probability that the investor makes money and receive a dividend, P( makes money and recceive dividend)  =  P(Get dividend |make money) * P(makes money)

= 0.43*0.28 = 0.1204 ~ 12%